Relationship between future value and present value long dash—Mixed stream. Using the information in the accompanying​ table, answer the questions that follow.

a. Determine the present value of the mixed stream of cash flows using a 55​% discount rate.

b. How much would you be willing to pay for an opportunity to buy this​ stream, assuming that you can at best earn 55​% on your​ investments?

c. What​ effect, if​ any, would a 77​% rather than a 55​% opportunity cost have on your​ analysis?

Year

​(t​)

Cash flow

1

​$700

2

​$1,000

3

​$1,000

4

​$1,600

5

​$1,900
Answer #1

a. Present Value = 700*PVF(55%,1 year) + 1,000*PVF(55%, 2 years) + 1,000*PVF(55%, 3 years) + 1,600*PVF(55%, 4 years) + 1,900*PVF(55%, 5 years)

= 700*0,645 + 1,000*0.416 1,000*0.269 + 1,600*0.173 + 1,900*0.112

= $1,626.1

b.Amount that should be paid = $1,626.1

c. A 77% return would decrease the present value of cash flows

= 700*0.565 + 1,000*0.319 + 1,000*0.180 + 1,600*0.102 + 1,900*0.058

= $1,167.9